OpenStudy (amy2012):

Solve the equation S/R(that's a fraction)+5=T/R(also a fraction) for R in terms of S and T

OpenStudy (utophii):

\( \frac{T-S} R = 5, \quad R = \frac{T-S} 5.\)

OpenStudy (sshayer):

correct.